OpenStudy (anonymous):

Write the ratio as a ratio of whole numbers using fractional notation. Write the fraction in simplest form. Remember to cancel units when applicable. 39 hours to 27 hours

OpenStudy (anonymous):

39/27 = 3 * 13/ 3 * 9 = 13/9